# DAIRstega: Dynamically Allocated Interval-Based Generative Linguistic Steganography with Roulette Wheel

### v4

@@ 5 Outlook
We have stimulated thinking in this field and discovered gaps in the field for scholars to conduct follow-up research.
1. Given the limited studies on longer and more diverse LS. The scheme can be designed to improve the length and diversity. 2. We know that fine-tuning LLMs brings huge costs, so the fine-tuning-free LS can be improved. Knowledge Graphs or Chain of Thought are used to enhance the prompts, and encoding text features to achieve a convenient and completely concealed LS scheme. 3. For watermarking, the idea of LS design can be used to take the hyperparameters of the open-source LLMs as the part that needs to be verified, and construct a new watermarking scheme. 4. Virtual social robots can be built, leveraging secrets to generate multi-turn conversations. It promises contactless covert communication, keeping the sender safe. 5. We recognize the shortcomings of current steganalysis for detecting DAIRstega’s stegos. So steganalysis needs to be developed.
Appendix A Supplementary
A.1 Related work
Perceptual concealment” focuses on ensuring that the scheme generates stego with complete and fluent. This is the most basic requirement of LS. In pursuit of this objective, Fang et al. (Fang, Jaggi, and Argyraki 2017) proposed an LSTM-based LS scheme. This scheme segments the vocabulary into several sets based on bit blocks. It then selects tokens with the highest probability that corresponds to the secret information from the candidate pool. Ding et al. (Ding et al. 2023a) combined the conditional generation strategy with the replacement technique, using text sequences as auxiliary data in the stego generation process to enhance the embedding capabilities. Yang et al. (Yang et al. 2019) trained a language model to generate stegos. Using HC-based (Huffman coding) and FLC-based (Fixed length coding) to encode the candidate pool, these stegos have excellent completeness and fluentness.
Statistical concealment” requires the stegos distribution to be close to the covers. To achieve this goal, Yang et al. (Yang et al. 2021) designed an encoder-decoder structure, and VLC-based (Variable length coding) is used to encode the CP. To further reduce the distribution difference between cover and stego, Zhang et al. (Zhang et al. 2021) used adaptive dynamic grouping coding to recursively embed secret. Zhou et al. (Zhou et al. 2021) used a GAN (Generative Adversarial Network) to design an adaptive probability distribution steganography. These ensure statistical concealment.
Semantic concealment” aims to generate the stego that is coherent and semantically controllable. To this end, Li et al. (Li et al. 2021) put forward an LS scheme based on the knowledge graph. Yang et al. (Yang et al. 2023) utilized semantic information encoding to embed secret information, realizing the effect of maintaining semantics and increasing the embedding capacity during the translation process. Wang et al. (Wang et al. 2023a) leveraged the relevance of social network context to enhance contextual semantic relevance while maintaining embedding rates. Lu et al. (Lu et al. 2023) used contextual learning and GPT2 guidance to generate stegos.
A.2 Supplement the effect of the proposed embedding way
Table 9: Text quality, statistical analysis, and discourse-matching comparison of stegos generated by DAIRstega and baselines. The embedding rates of each way include 1 Bpw (bit per word) and 2 Bpw. “Avg.” represents the overall performance. Bold is the best overall result. “ *” is the suboptimal overall result. The meanings of “bit” are the same in Table 2.
Embedding Param / Bpw Text quality Statistical analysis Discourse-matching
PPL $\downarrow$ $\Delta$Pcs $\downarrow$ CS $\uparrow$ JSD $\downarrow$ ED $\downarrow$ MD $\downarrow$ $\Delta$DP $\downarrow$ $\Delta$LDA $\downarrow$ Mauve $\uparrow$ BLEU $\uparrow$ Score $\uparrow$
LLMs+HC bit = 1 / 1.00 13.331 5.639 96.54 45.79 0.263 8.255 3.46 0.014 70.60$\pm$27.36 54.26$\pm$4.25 58.44$\pm$4.07
bit = 2 / 2.52 24.334 16.641 90.57 47.35 0.350 9.354 9.43 0.009 64.88$\pm$30.03 55.00$\pm$5.77 56.19$\pm$4.01
Avg. 18.832 11.140 93.55 46.57 0.307 8.805 6.45 0.011 67.74 54.63 57.32
LLMs+ADG $\tau$ = 0.3 / 1.34 12.750 5.058 97.20 45.59 0.240 8.316 2.80 0.010 77.86$\pm$23.65 71.79$\pm$3.03 59.18$\pm$5.84
$\tau$ = 0.65 / 2.99 16.175 8.483 95.95 44.74 0.331 8.478 4.05 0.018 75.60$\pm$27.02 62.60$\pm$4.68 58.60$\pm$5.69
Avg. 14.463* 6.770* 96.58 45.17* 0.285 8.397* 3.42 0.014 76.73* 67.20* 58.35*
LLMs+AC CP = 8 / 1.25 73.454 65.762 82.16 51.64 0.746 15.415 17.84 0.014 35.64$\pm$22.65 50.92$\pm$29.30 45.10$\pm$3.07
CP = 32 / 1.32 86.238 78.545 80.38 51.21 0.829 14.194 19.62 0.012 37.18$\pm$24.59 48.01$\pm$27.45 47.51$\pm$2.76
Avg. 79.846 72.154 81.27 51.43 0.788 14.805 18.73 0.013* 36.41 49.47 46.31
LLMs+FLC bit = 1 / 1.34 12.883 5.190 97.04 46.56 0.241 8.280 2.96 0.031 79.48$\pm$26.53 54.88$\pm$6.33 58.44$\pm$4.64
bit = 2 / 2.67 30.481 22.788 93.86 46.30 0.328 8.939 6.14 0.011 61.10$\pm$29.30 58.21$\pm$9.88 55.20$\pm$3.39
Avg. 21.682 13.989 95.45 46.43 0.284* 8.610 4.55 0.021 70.29 56.55 56.82
Ours $\beta$=1 $\alpha$=8 / 1.10 7.686 0.006 97.20 38.86 0.237 6.458 2.80 0.017 79.91$\pm$26.75 77.88$\pm$30.40 65.76$\pm$4.13
$\beta$=1 $\alpha$=32 / 1.11 8.184 0.492 97.65 38.48 0.217 6.293 2.35 0.022 76.97$\pm$26.47 77.85$\pm$30.42 65.22$\pm$4.61
$\beta$=1 $\alpha$=48 / 1.13 8.219 0.527 97.45 38.46 0.226 6.273 2.55 0.020 78.70$\pm$25.51 76.53$\pm$31.71 65.25$\pm$4.78
Avg. (1Bpw) 8.030 0.342 97.43 38.60 0.227 6.341 2.57 0.020 78.53 77.42 65.41
[…]
$\beta$=0.5 $\alpha$=32 / 2.56 20.876 13.184 92.43 39.69 0.389 8.976 7.57 0.029 75.58$\pm$20.11 66.99$\pm$27.08 60.89$\pm$3.85
$\beta$=0.5 $\alpha$=48 / 2.58 22.069 14.377 92.44 38.31 0.389 8.592 7.56 0.006 72.40$\pm$26.33 79.64$\pm$12.47 60.29$\pm$4.00
Avg. (2Bpw) 18.907 11.214 94.09 38.60 0.336 8.258 5.91 0.018 76.27 73.27 61.22
[…]
Table 10: Anti-steganalysis comparison of stegos generated by DAIRstega and baselines. The embedding rates of each way include 1Bpw (bit per word) and 2Bpw. “Avg.” represents the overall performance. Bold is the best overall result. “ *” is the suboptimal overall result.
Embedding Param / Bpw LS_CNN TS_CSW EILG UP4LS
LLMs+HC bit = 1 / 1.00 77.62$\pm$1.38 78.41$\pm$1.65 78.09$\pm$1.33 78.25$\pm$1.21 78.58$\pm$1.69 78.29$\pm$2.24 86.73$\pm$1.32 86.13$\pm$0.50
bit = 2 / 2.52 77.54$\pm$1.50 77.57$\pm$2.05 76.03$\pm$1.74 76.09$\pm$1.33 78.29$\pm$0.72 78.95$\pm$1.47 88.19$\pm$1.00 88.43$\pm$2.61
Avg. 77.58 77.99 77.06 77.17 78.44 78.62 87.46 87.28
LLMs+ADG $\tau$ = 0.3 / 1.34 70.10$\pm$0.66 70.17$\pm$0.89 69.45$\pm$2.05 69.34$\pm$2.53 70.99$\pm$3.15 70.49$\pm$1.11 84.00$\pm$0.83 84.23$\pm$1.07
$\tau$ = 0.65 / 2.99 63.06$\pm$2.49 63.69$\pm$3.61 62.22$\pm$1.46 61.86$\pm$1.18 66.14$\pm$1.38 66.15$\pm$1.38 80.42$\pm$0.59 80.57$\pm$0.58
Avg. 66.58* 66.93* 65.84* 65.60* 68.57* 68.32* 82.21* 82.40*
LLMs+AC CP = 8 / 1.25 90.15$\pm$1.17 89.98$\pm$1.21 89.69$\pm$0.93 89.37$\pm$0.88 86.92$\pm$1.12 86.52$\pm$0.96 93.62$\pm$0.89 93.37$\pm$0.40
CP = 32 / 1.32 88.45$\pm$1.38 88.44$\pm$1.25 90.80$\pm$1.97 90.52$\pm$2.18 86.44$\pm$2.47 85.86$\pm$3.10 94.03$\pm$0.64 94.16$\pm$0.95
Avg. 89.30 89.21 90.25 89.95 86.68 86.19 93.83 93.77
LLMs+FLC bit = 1 / 1.34 78.53$\pm$1.20 80.05$\pm$1.39 82.72$\pm$1.27 83.32$\pm$1.10 76.67$\pm$0.85 76.15$\pm$1.96 89.03$\pm$0.25 89.20$\pm$0.80
bit = 2 / 2.67 76.49$\pm$1.15 76.54$\pm$1.13 77.08$\pm$2.06 77.26$\pm$1.75 76.95$\pm$1.69 76.46$\pm$1.74 90.62$\pm$0.89 90.37$\pm$0.40
Avg. 77.51 78.30 79.90 80.29 76.81 76.31 89.83 89.79
Ours $\beta$=1 $\alpha$=8 / 1.10 59.50$\pm$2.77 60.15$\pm$10.59 56.82$\pm$4.09 64.21$\pm$10.03 67.25$\pm$3.23 68.04$\pm$1.53 75.40$\pm$0.37 76.26$\pm$1.28
$\beta$=1 $\alpha$=32 / 1.11 59.35$\pm$1.25 59.99$\pm$5.51 54.05$\pm$4.72 61.68$\pm$11.17 62.28$\pm$0.50 62.66$\pm$2.31 71.97$\pm$0.64 72.75$\pm$1.21
$\beta$=1 $\alpha$=48 / 1.13 58.00$\pm$2.37 56.08$\pm$13.79 54.98$\pm$3.14 59.17$\pm$11.37 65.51$\pm$2.23 66.78$\pm$3.91 71.40$\pm$0.65 70.89$\pm$2.05
Avg. (1Bpw) 58.95 58.74 55.28 61.69 65.01 65.83 72.92 73.30
[…]
$\beta$=0.5 $\alpha$=32 / 2.56 72.06$\pm$1.17 71.78$\pm$1.21 68.73$\pm$1.29 68.19$\pm$3.44 71.71$\pm$1.95 70.34$\pm$2.41 76.41$\pm$2.56 75.90$\pm$2.73
$\beta$=0.5 $\alpha$=48 / 2.58 70.17$\pm$1.81 70.40$\pm$3.11 69.48$\pm$0.68 67.80$\pm$1.98 70.64$\pm$1.61 71.00$\pm$2.23 79.46$\pm$0.56 78.28$\pm$1.13
Avg. (2Bpw) 68.37 68.13 67.11 65.25 69.34 69.06 74.97 74.14
[…]
A.3 Complete data for Table 4 and Table 5
In this section, we give the complete data of Table 4 and Table 5, as shown in Table 11 and Table 12. Table 11 contains the standard deviations of many datasets with different discourse, and Table 12 contains the standard deviations of multiple runs.
Table 11: Complete data for Table 4. The meanings of “bin” and “bit” are the same in Table 2. “a±$\pm$±b” represents “Avg.$\pm$±Std”. Bold is the best result. “ *” is the suboptimal result. The “Std”s of DAIRstega Mauve and BLEU values are bigger. This is not due to the instability of 5-time runs, but is obtained from the effect on dozens of different discourse datasets, and represents the overall performance.
Schemes Param / Bpw Mauve $\uparrow$ BLEU $\uparrow$ Score $\uparrow$
Fang bin = 1 / 1.00 2.15$\pm$0.87 0.90$\pm$1.58 42.58$\pm$1.30
bin = 3 / 3.00 1.96$\pm$0.68 0.90$\pm$1.59 41.65$\pm$1.33
RNNstega bit = 1 / 1.00 12.73$\pm$10.56 1.87$\pm$3.28 52.57$\pm$2.20
bit = 3 / 2.63 11.97$\pm$10.56 1.83$\pm$3.21 52.36$\pm$1.80
ADG $\tau$ = 0.5 / 4.38 18.54$\pm$13.89 2.72$\pm$4.79 47.55$\pm$2.25
LSCS - / 1.12 2.40$\pm$1.78 31.95$\pm$12.89 59.22$\pm$3.14
PLMmark - / - 7.22$\pm$4.88 0.15$\pm$0.27 43.50$\pm$0.79
DAIRstega $\beta$ = 1 $\alpha$ = 8 / 1.10 79.91$\pm$26.75 76.37$\pm$31.85 65.76$\pm$4.13
$\alpha$ = 32 / 1.11 76.97$\pm$26.47 77.85$\pm$30.42* 65.22$\pm$4.61
$\alpha$ = 48 / 1.13 78.70$\pm$25.51 76.53$\pm$31.71 65.25$\pm$4.78*
[…]
$\alpha$ = 32 / 2.56 75.58$\pm$20.11 66.99$\pm$27.08 60.89$\pm$3.85
$\alpha$ = 48 / 2.58 72.40$\pm$26.33 79.64$\pm$12.47 60.29$\pm$4.00
Table 12: Complete data for Table 5. The meanings of “bin” and “bit”are the same in Table 2. “a±$\pm$±b” represents “Avg.$\pm$±Std”. “Avg” and “Std” in this table are obtained by running 5 times. Bold is the best result. “ *” is the suboptimal result.
Schemes Param / Bpw LS_CNN TS_CSW EILG UP4LS
Fang bin = 1 / 1.00 99.70$\pm$0.29 99.68$\pm$0.31 88.34$\pm$21.22 91.01$\pm$15.69 99.50$\pm$0.25 99.51$\pm$0.25 99.75$\pm$0.06 99.60$\pm$0.09
bin = 3 / 3.00 99.75$\pm$0.16 99.76$\pm$0.15 89.03$\pm$20.58 92.39$\pm$13.82 99.59$\pm$0.17 99.52$\pm$0.66 99.88$\pm$0.10 99.80$\pm$0.15
RNNstega bit = 1 / 1.00 85.81$\pm$1.93 86.03$\pm$1.40 81.39$\pm$1.70 82.59$\pm$1.62 85.86$\pm$0.50 86.29$\pm$0.89 97.42$\pm$0.44 95.97$\pm$0.67
bit = 3 / 2.63 82.93$\pm$2.18 84.48$\pm$2.20 80.35$\pm$2.38 77.94$\pm$3.04 82.71$\pm$0.17 83.10$\pm$1.02 96.78$\pm$0.82 94.74$\pm$1.34
ADG $\tau$ = 0.5 / 4.38 85.86$\pm$1.73 86.35$\pm$1.82 82.23$\pm$2.40 82.08$\pm$2.96 84.86$\pm$1.74 84.10$\pm$1.80 98.34$\pm$0.39 97.41$\pm$0.58
LSCS - / 1.12 78.86$\pm$3.07 78.98$\pm$2.83 81.19$\pm$6.62 80.11$\pm$8.26 84.15$\pm$0.41 85.11$\pm$0.79 96.22$\pm$2.08 95.48$\pm$3.11
PLMmark - / - 98.64$\pm$0.92 98.60$\pm$0.95 98.81$\pm$1.83 98.75$\pm$1.93 99.92$\pm$0.17 99.92$\pm$0.11 99.85$\pm$0.30 99.85$\pm$0.30
DAIRstega $\beta$ = 1 $\alpha$ = 8 / 1.10 59.50$\pm$2.77 60.15$\pm$10.59 56.82$\pm$4.09 64.21$\pm$10.03 67.25$\pm$3.23 68.04$\pm$1.53 75.40$\pm$0.37 76.26$\pm$1.28
$\alpha$ = 32 / 1.11 59.35$\pm$1.25 59.99$\pm$5.51* 54.05$\pm$4.72 61.68$\pm$11.17 62.28$\pm$0.50 62.66$\pm$2.31* 71.97$\pm$0.64 72.75$\pm$1.21
$\alpha$ = 48 / 1.13 58.00$\pm$2.37 56.08$\pm$13.79 54.98$\pm$3.14* 59.17$\pm$11.37 65.51$\pm$2.23 66.78$\pm$3.91 71.40$\pm$0.65* 70.89$\pm$2.05*
[…]
$\alpha$ = 32 / 2.56 72.06$\pm$1.17 71.78$\pm$1.21 68.73$\pm$1.29 68.19$\pm$3.44 71.71$\pm$1.95 70.34$\pm$2.41 76.41$\pm$2.56 75.90$\pm$2.73
$\alpha$ = 48 / 2.58 70.17$\pm$1.81 70.40$\pm$3.11 69.48$\pm$0.68 67.80$\pm$1.98 70.64$\pm$1.61 71.00$\pm$2.23 79.46$\pm$0.56 78.28$\pm$1.13
A.4 Supplement data for anti-steganalysis
We also perform other 7 steganalysis methods to detect the stegos we generated, as shown in Table A.4. These include non-BERT-based, BERT-based, and LLMs-based steganalysis methods.
Table 13: The supplemental anti-steganalysis capability of the DAIRstega from other 7 steganalysis works. Bold represents the best result and the best Avg. result. “ *” represents the suboptimal result.
@@ A.4 Supplement data for anti-steganalysis
  • [1] VAE-Stega: Linguistic Steganography Based on Variational Auto-Encoder. IEEE Trans. Inf. Forensics and Security, 880–895, 2021.
  • [2] A Fast and Efficient Text Steganalysis Method. IEEE Signal Processing Lett., 26(4): 627–631, 2019.
  • [3] A Hybrid R-BiLSTM-C Neural Network Based Text Steganalysis. IEEE Signal Processing Lett., 26(12): 1907–1911, 2019.
  • [4] High-performance Linguistic Steganalysis, Capacity Estimation and Steganographic Positioning. In Proceeding of IWDW, 2021.
  • [5] Small-Scale Linguistic Steganalysis for Multi-Concealed Scenarios. IEEE Signal Processing Lett., 29: 130–134, 2022.
  • [6] SeSy: Linguistic Steganalysis Framework Integrating Semantic and Syntactic Features. IEEE Signal Processing Lett., 31–35, 2021.
  • [7] Towards Next-Generation Steganalysis: LLMs Unleash the Power of Detecting Steganography. arXiv:2405.09090, 2024.
  • [8] Linguistic Steganalysis via LLMs: Two Modes for Efficient Detection of Strongly Concealed Stego. arXiv:2406.04218, 2024.
A.5 Example of cross-modal secret information hiding
Table 14: Examples of longer stegos generated by DAIRstega using cross-modal information. Bitstream is the binary corresponding to secret.

### v5

@@ 6 Outlook
By reflecting on this field, we found gaps in it that can be studied by scholars. 1. This paper preliminarily explores the performance of the proposed method in long stegos. However, the research on long stegos is still in the basic stage. So, given the limited studies on longer and more diverse steganography, the scheme can be designed to improve the length and diversity. 2. There is currently a lack of high-quality steganography schemes for closed-source LLMs. So, knowledge Graphs or Chain of Thought can enhance the prompts, while encoding text features allows for convenient and completely concealed steganography. 3. The scheme in this paper can also be used as a watermarking for LLMs, which can make up for the insufficiency of the watermarking for LLMs being difficult to completely extract. Therefore, a new watermarking for LLMs can be designed by referring to the scheme in this paper. 4. At present, linguistic steganography can achieve strong concealment and anti-steganalysis capabilities, that is, it is relatively safe. However, the sender of the stego is dangerous. Therefore, virtual social robots can be developed that utilize secrets to generate multi-turn conversations, promising contactless, covert communication that ensures the sender’s safety. 5. Finally, we found that existing steganalysis methods have poor detection performance on DAIRstega’s stegos. Therefore, it is essential to advance steganalysis techniques.
Acknowledgment
We would like to express our gratitude to the MindSpore Community for their invaluable support and resources throughout this research. Thanks for the support provided by MindSpore Community. Thank the editor and reviewers of “Applied Soft Computing” for their review and constructive suggestions.
Declaration of competing interest
The authors declare no conflict of interest.
Funding
This work is supported by the National Natural Science Foundation of China (Grant U21B2020) and supported by BUPT Excellent Ph.D. Students Foundation (Grant CX2023120, CX20241055).
Author contributions
Yihao Wang: Conceptualization; Data curation; Investigation; Methodology; Software; Funding acquisition; original draft. Ruiqi Song: Data curation; Investigation; Resources; Software; Validation; original draft. Lingxiao Li: Data curation; Resources; Software; Funding acquisition; original draft; review & editing. Yihao Wang and Lingxiao Li contributed equally to this work and should be considered co-first authors. Ru Zhang: Funding acquisition; Investigation; Project administration; Supervision; review & editing. Jianyi Liu: Formal analysis; Supervision; Investigation; Validation; review & editing.
Ethical Statement
This research does not involve human participants, personal data, or animal subjects; therefore, it does not raise any direct ethical concerns. The data used in this study were sourced from publicly available datasets or generated by applicable laws and regulations.
@@ Acknowledgment
7.4 Example of cross-modal secret information hiding
Table 15: Examples of longer stegos generated by DAIRstega using cross-modal information. Bitstream is the binary corresponding to the secret.